OpenStudy (anonymous):

Evaluate the limit. limx→0{(3.5+4.5x)^3−42.875}/x=

hartnn (hartnn):

just plug in x=0

OpenStudy (anonymous):

I tried, doesn't it just equal 0/0?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

-42.875/0 which is \(-\infty \)

hartnn (hartnn):

oops sorry, i misread the question

hartnn (hartnn):

ok, so 3.5^3 = 42.875

hartnn (hartnn):

we have the form of \(a^3-b^3\) here right ?

OpenStudy (anonymous):

@hartnn I tried all these already and nothing seems to be working. I've working on this problem for half an hour now, thats why I'm asking on here finally

hartnn (hartnn):

use the formula \((a-b)(a^2+ab+b^2)\) a= 3.5 +4.5 x b= 3.5

OpenStudy (anonymous):

@hartnn so after plugging in the values for a &b I should have the answer?

hartnn (hartnn):

did you try plugging in? after that there will be 'x' in the numerator which will get cancelled with the 'x' in the denominator and then you can directly plug in x=0

OpenStudy (anonymous):

so I got : (3.5 - 4.5x)(3.5^2+15.75x+3.5^2)

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

is that the correct equation?

hartnn (hartnn):

a-b is just 4.5 x isn't it ?

hartnn (hartnn):

you took a = 3.5, b= 4.5x thats not correct

hartnn (hartnn):

a= 3.5 +4.5x b=3.5

OpenStudy (anonymous):

ok so then the correct equation is: (4.5x)(3.5^2+15.75x+3.5^2) ?

OpenStudy (anonymous):

or (4.5x)(7^2+15.75x)? @hartnn

hartnn (hartnn):

no...a= 3.5 +4.5x so a^2 =(3.5 +4.5x)^2

OpenStudy (anonymous):

I'm kind of confused now. so if I calculated your equation the answer would be = 23.75? @hartnn

OpenStudy (anonymous):

@hartnn nevermind that answer didn't work either

hartnn (hartnn):

show each of your steps, i'll help u spot the error

OpenStudy (anonymous):

okay hold on let me type it out

OpenStudy (anonymous):

Formula: (a−b)(a2+ab+b2) a= 3.5 +4.5 x b= 3.5 3.5+4.5x - 3.5 = 4.5x (4.5x)({[3.5+4.5x]*2}+{[3.5+4.5x]*3.5}+[3.5*2])

OpenStudy (anonymous):

@hartnn that's the equation I got

hartnn (hartnn):

ok, thats correct now cancel out x and plug in x=0 in the remaining part

hartnn (hartnn):

the last part is 3.5^2 not 3.5*2

hartnn (hartnn):

\((4.5x)({[3.5+4.5x]^2}+{[3.5+4.5x]*3.5}+[3.5^2])\)

hartnn (hartnn):

\((4.5)({[3.5+4.5(0)]^2}+{[3.5+4.5(0)]*3.5}+[3.5^2])\)

OpenStudy (anonymous):

@hartnn I might be wrong but so is 70.875 the answer?

hartnn (hartnn):

using calculator, its 165.375

hartnn (hartnn):

4.5*3*3.5^2

OpenStudy (anonymous):

Yeah that was the correct answer, but how did you get the 4.5*3*3.5^2? I thought the first two parts of the equation equaled 0? so it would be 4.5*3.5+3.5^2

hartnn (hartnn):

\((4.5)({[3.5+4.5(0)]^2}+{[3.5+4.5(0)]*3.5}+[3.5^2])\) \((4.5)({[3.5]^2}+{[3.5]*3.5}+[3.5^2]) \\ 4.5 ([3.5^2]+[3.5^2]+[3.5^2])\)

hartnn (hartnn):

a+a+a=3a

OpenStudy (anonymous):

@hartnn oh okay, I see now!

OpenStudy (anonymous):

@hartnn thank you so much

hartnn (hartnn):

welcome ^_^